OpenStudy (anonymous):

whats the answer for x-4y=-4; x=-2,4,6

OpenStudy (anonymous):

\[2-4y=-4 \]\[-4y=-6\]\[y=1.5\] \[4-4y=-4\]\[-4y=-8\]\[y=2\] \[6-4y=-4\]\[-4y=-10\]\[y=2.5\]

OpenStudy (anonymous):

Oops, made a mistake on the first one: \[-2-4y=-4\]\[-4y=-2\]\[y=0.5\]

OpenStudy (anonymous):

Oh I got it thanks I was stuck on it thanks:-)